OpenStudy (anonymous):

i need help i dont know ratios The numerator and denominator of a fraction are in the ratio of 3 to 5. If the numerator and denominator are both increased by 2, the fraction is now equal to 2/3. If n = the numerator and d = the denominator, which of the following systems of equations could be used to solve the problem? 5n = 3d and 3n + 6 = 2d + 4 5n = 3d and 4n + 4 = 3d + 6 3n = 5d and 3n + 6 = 2d + 4

hartnn (hartnn):

the fraction is now equal to what ?

OpenStudy (anonymous):

im sorry ill fix it

hartnn (hartnn):

ok, then ratio = numerator/denominator = n/d = 3/5 so, can you simplify n/d=3/5 ? what u get ? (cross-multiply...)

OpenStudy (anonymous):

ill try

OpenStudy (anonymous):

3d/5n

hartnn (hartnn):

in fact 5n=3d now, the 2nd part says both n and d are increased by 2,so n--->n+2 d--->d+2 (n+2)/(d+2) =2/3 again cross multiply!

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

i git it now